i need help with the question 13.2


13.2 Answer the following for the data of Exercise 12.11: a. What is the 10-percentile and 90-percentile from the sample? b. What is the mean number of disk VOs per program? c, what is the 90% confidence interval for the mean? d. What fraction of programs make less than or equal to 25 LOs and what is the 90% confidence interval for the fraction? e, what is the one-sided 90% confidence interval for the mean?
media%2Ffd1%2Ffd14c7c5-50b6-4bab-88c0-45
Answer #1

Following is the ordered data set and calculations for mean and SD:

228 6691727 1 059222244 1137 20336-1. 9 2 4 5 5 5 5 8 8 3 1 1 4 6770 55222677 2 2596441631111 3 2 11238 9 ! 3 4 5 5 6 9 1 3 3 3 3 4 4 5 8 8 9 1 3 3 4 4 4 5 5 6689258 111111122222222222333333333333448 8

Sample size: n=33

(a)

The 10th percentile will be

P_{10}=rac{10}{100}cdot (33+1)=3.4th, data, value

Since 3rd data value is 13 and 4th data value is 14 so required percentile is

P_{10}=13+(14-13)cdot 0.4=13.4

--------------------

The 90th percentile will be

90 100 P9o (33 +1)-30.6th data val ue

Since 30th data value is 38 and 31st data value is 39 so required percentile is

P_{90}=38+(39-38)cdot 0.60=38.60

b)

Mean:

ar{x}=rac{sum x}{n}=rac{888}{33}=26.91

c)

Standard deviation:

s=sqrt{rac{sum left ( x-ar{x} ight )^{2}}{n-1}}=sqrt{rac{2884.7273}{33-1}}approx 9.4946

Confidence interval:

d)

Out of 33 data values, 16 are less than equal to 25 so required fraction is

25 p = 0.758

Confidence interval:

e)

The critical value for one sided confidence interval is z_{0.90}=1.28

The one sided confidence interval for mean is

·24336 9.4946 = 29.08 +tcriticalyM = 26.96 + 1.28 Vn

The one sided confidence interval is (-infinity, 29.08).